OpenStudy (anonymous):

If you have 430.0 mL of water at 25.00 °C and add 140.0 mL of water at 95.00 °C, what is the final temperature of the mixture? Use 1.00 g/mL as the density of water.

OpenStudy (anonymous):

I just need help with an equation to use. i've looked over all my notes and can't seem to come up with anything. I don't really want an answer, more of a bit of help.

OpenStudy (anonymous):

you will also need heat capacity of water... and then you go with this equation: Q = m * cp * (T2-T1) where Q = heat m = mass cp = heat capacity of water (T2-T1) = temperature now since you have heating of colder water and cooling of hotter water you write: Q(hot) = - Q(cold) and T2 is your final temperature on both sides so you have: m(h) * cp(w) * (T2 - 95) = m(c) * cp(w) * (T2 - 25) where indexes are: h - hot c - cold w - water and all you need to do is to solve that equation for T2 and you will get your answer! hope this helps ;)